OpenStudy (whystudyalone3):

If a dose = 150 mg, how many doses are in 3 g?

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (skullpatrol):

Recall 1 g = 1,000 mg

OpenStudy (whystudyalone3):

I think its 20 doses

OpenStudy (whystudyalone3):

I got it right. Do you think that you can help me with one that I've been stuck on for a while now?

OpenStudy (skullpatrol):

I can try :-)

OpenStudy (whystudyalone3):

Heres the question: One number is twice as large as the second. The sum of the two numbers is 366. What is the smaller number?

OpenStudy (unklerhaukus):

n + 2n = 366 solve for n

OpenStudy (skullpatrol):

Let x = smaller number. Then 2x = larger number.

OpenStudy (skullpatrol):

smaller + larger = 366 x + 2x = 366

OpenStudy (whystudyalone3):

I'm still not really sure what to do... Can someone please walk me through it?

OpenStudy (skullpatrol):

What is x + 2x = ?

OpenStudy (whystudyalone3):

x + 2x = 366

OpenStudy (skullpatrol):

x + 2x = ( 1 + 2 ) x = 3x

OpenStudy (whystudyalone3):

I'm not sure how to solve that equation ;(

OpenStudy (skullpatrol):

Replace x + 2x with 3x

OpenStudy (whystudyalone3):

ohh okay so now it would be 3x = 366? Then divide 366 by 3?

OpenStudy (unklerhaukus):

yes

OpenStudy (skullpatrol):

Yes, that means you are dividing both sides by 3.

OpenStudy (whystudyalone3):

okay, so the answer I got was 122. Is that right?

OpenStudy (skullpatrol):

Correct.

OpenStudy (mathmale):

Find both numbers, and then check your solution through substituting them back into the original problem (statement). Try to make a habit of checking everything you can.

OpenStudy (mathmale):

What is the larger number?

OpenStudy (whystudyalone3):

Thank you all so much!

OpenStudy (skullpatrol):

Thanks for trying to learn :-)